# Critical point representation of the mutual information in the sparse stochastic block model

### v1

@@ 4 Cavity computation
[…]
@@ 5 Critical point representation
To relate the Gateaux derivative of the limit free energy to that of the initial condition, it will be convenient to recall from [18, Remark 3.2] that for any $\mu\in\mathcal{M}_{+}$, the density of the Gateaux derivative $D_{\mu}\psi(\mu)$ is
$D_{\mu}\psi(\mu,x)=\operatorname{\mathbb{E}}\langle c+\Delta\epsilon x\rangle_{*}\log\langle c+\Delta\epsilon x\rangle_{*}-c-\Delta\overline{m}x,$ (5.1)
@@
$H^{\mu}(\epsilon):=-\int_{-1}^{1}(c+\Delta\epsilon x)\,\mathrm{d}\mu(x)+\sum_{x\in\Pi_{*}(\mu)}\log(c+\Delta\epsilon x).$ (5.2)
Here $\Pi_{*}(\mu)$ denotes a Poisson point process on $[-1,1]$ with mean measure $(c+\Delta\sigma^{*}x)\,\mathrm{d}\mu(x)$. This means that for any bounded and measurable function $f=f(\epsilon^{1},\ldots,\epsilon^{n})$ of finitely many replicas,
$\langle f(\epsilon^{1},\ldots,\epsilon^{n})\rangle_{*}:=\langle f\rangle_{*}:=\frac{\int_{\Sigma_{1}^{n}}f(\epsilon^{1},\ldots,\epsilon^{n})\prod_{\ell\leq n}\exp H^{\mu}(\epsilon^{\ell})\,\mathrm{d}P^{*}(\epsilon^{\ell})}{\big{(}\int_{\Sigma_{1}}\exp H^{\mu}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)\big{)}^{n}}.$ (5.3)
It will also be convenient to remember from the proof of [18, Lemma 2.4] that the Gateaux derivative of the enriched free energy (1.22) at $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ in the direction of a probability measure $\eta\in\Pr[-1,1]$ is given by
@@
[…]
@@
It follows that $\nu$ and $\nu^{\prime}$ have all the same moments of order $n\geq 2$, and are therefore equal by the Stone-Weierstrass theorem. ∎
@@
$-c-\frac{\Delta\operatorname{\mathbb{E}}w_{1}}{N+1}\sum_{i\leq N+1}\sigma_{i}+\operatorname{\operatorname{\mathcal{O}}}(N^{-1})=-c-\Delta\overline{m}\operatorname{\mathbb{E}}w_{1}+\operatorname{\operatorname{\mathcal{O}}}(N^{-1}).$
@@
$D_{\mu}\overline{F}_{N+1}(t,\mu,\lambda^{N};\eta)=(N+1)\operatorname{\mathbb{E}}\log\big{\langle}(c+\Delta\epsilon w_{1})^{\widetilde{G}_{1}^{w}}\big{\rangle}_{N+1}-c-\Delta\overline{m}\operatorname{\mathbb{E}}w_{1}.$ (5.10)
@@
$D^{1}$ $:=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\big{\langle}(c+\Delta\epsilon w_{1})^{\widetilde{G}_{1}^{w}}\exp z_{N}^{t,\mu}(\epsilon,\sigma)\big{\rangle}^{\prime}-\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}^{t,\mu}(\epsilon,\sigma)\rangle^{\prime}\,\mathrm{d}P^{*}(\epsilon)$
@@
$D^{1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}e^{-c(t+s)-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}$ $(c+\Delta\epsilon w_{1})^{{G^{\prime}}^{w}}\exp z_{N}^{\prime}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)$
$-\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}e^{-c(t+s)-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}\exp z_{N}^{\prime}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)+o_{N}(1),$
@@
$\mathbb{P}\big{\{}{G^{\prime}}^{w}=1\mid\epsilon^{*},w_{1}\big{\}}:=\frac{c+\Delta\epsilon^{*}w_{1}}{N+1}.$
@@
$D^{1}=\operatorname{\mathbb{E}}\log\big{\langle}(c+\Delta\epsilon w_{1})^{{G^{\prime}}^{w}}\big{\rangle}_{*}+o_{N}(1)=\frac{1}{N+1}\operatorname{\mathbb{E}}\langle c+\Delta\epsilon w_{1}\rangle_{*}\log\langle c+\Delta\epsilon w_{1}\rangle_{*}+o_{N}(1).$
@@
$D_{\mu}f(t,\mu;\eta)=\operatorname{\mathbb{E}}\langle c+\Delta\epsilon w_{1}\rangle_{*}\log\langle c+\Delta\epsilon w_{1}\rangle_{*}-c-\Delta\overline{m}\operatorname{\mathbb{E}}w_{1}.$
@@
$D_{\mu}\psi(\mu+t\nu,x)=\operatorname{\mathbb{E}}g(\langle\epsilon\rangle_{*}x)=(c+\Delta\operatorname{\mathbb{E}}\langle\epsilon\rangle_{*}x)\log(c)+c\sum_{n\geq 2}\frac{(-\Delta/c)^{n}}{n(n-1)}\operatorname{\mathbb{E}}\langle\epsilon\rangle_{*}^{n}x^{n}-c.$ (5.12)
The equation $G_{\nu}=D_{\mu}\psi(\mu+t\nu,\cdot)$ is therefore equivalent to the equality between all moments of $y_{1}$ and $\langle\epsilon\rangle_{*}$ of order $2$ and higher. By the Stone-Weierstrass theorem, this is in turn equivalent to the equality between the laws of $y_{1}$ and $\langle\epsilon\rangle_{*}$. These laws are equal if and only if the fixed point equation $\nu=\Gamma_{t,\mu}(\nu)$ is satisfied. This completes the proof. ∎
@@ 6 The disassortative and small-time settings
in such a way that $\varphi(1)=\overline{F}_{N}(t,\mu)$ and $\varphi(0)=\psi(\mu+t\nu)$. We denote by $x_{1}$ a random variable sampled from $\nu$. Recall from the derivative computations in [18, Corollaries 2.2 and 2.5] that
@@
    $W\big{(}\Gamma_{t,\mu}(\nu),\Gamma_{t^{\prime},\mu^{\prime}}(\nu)\big{)}\leq C\big{(}\lvert t^{\prime}-t\rvert+\mu[-1,1]W\big{(}\overline{\mu},\overline{\mu}^{\prime}\big{)}+\lvert\mu[-1,1]-\mu^{\prime}[-1,1]\rvert\big{)}.$ (6.4)
@@
    Fix $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ as well as $\nu,\eta\in\mathcal{M}_{p}$. Conditionally on the ground truth signal $\epsilon^{*}$ sampled from $P^{*}$, introduce the measures
    $\,\mathrm{d}\tilde{\nu}(x):=(c+\Delta\epsilon^{*}x)\,\mathrm{d}\nu(x)\quad\text{and}\quad\,\mathrm{d}\tilde{\eta}(x):=(c+\Delta\epsilon^{*}x)\,\mathrm{d}\eta(x).$ (6.5)
    We denote by $(X_{k})_{k\geq 1}$ and $(Y_{k})_{k\geq 1}$ i.i.d. samples from the probability measures $\nu^{\prime}$ and $\eta^{\prime}$ induced by $\tilde{\nu}$ and $\tilde{\eta}$, and let $\Pi_{t}$ be a Poisson random variable with mean $t(c+\Delta\epsilon^{*}\overline{m})$. For each $u\in[0,1]$, we define the interpolating Hamiltonian $H_{u}$ on $\Sigma_{1}$ by
    $H_{u}(\epsilon):=-\epsilon\overline{m}+\sum_{x\in\Pi_{0,\mu}(0)}\log(c+\Delta\epsilon x)+\sum_{k\leq\Pi_{t}}\log\big{(}c+\Delta\epsilon(X_{k}+u(Y_{k}-X_{k}))\big{)}$
@@
    $Z_{u}:=\frac{\int_{\Sigma_{1}}\epsilon\exp H_{u}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)}{\int_{\Sigma_{1}}\exp H_{u}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)}.$
@@
    $\frac{\,\mathrm{d}}{\,\mathrm{d}u}Z_{u}$ $=\sum_{k\leq\Pi_{t}}(Y_{k}-X_{k})\bigg{(}\bigg{\langle}\frac{\epsilon}{c+\Delta\epsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg{\rangle}_{u}-\langle\epsilon\rangle_{u}\bigg{\langle}\frac{1}{c+\Delta\epsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg{\rangle}_{u}\bigg{)}.$
@@
    $F_{\nu^{\prime}}(x)=\frac{1}{c+\Delta\epsilon^{*}\overline{m}}\bigg{(}cF_{\nu}(x)+\Delta\epsilon^{*}\Big{(}xF_{\nu}(x)-\int_{-1}^{x}F_{\nu}(z)\,\mathrm{d}z\Big{)}\bigg{)}.$
@@
    $H_{u}(\epsilon):=-\epsilon\overline{m}$ $+\sum_{x\in\Lambda(t\tilde{\nu})}\log(c+\Delta\epsilon x)+u\sum_{x\in\Lambda((t^{\prime}-t)\tilde{\nu})}\log(c+\Delta\epsilon x)$
    $+\sum_{k\leq\Pi_{\tilde{s}}}\log\big{(}c+\Delta\epsilon(X_{k}+u(Y_{k}-X_{k}))\big{)}+u\sum_{k\leq\Pi_{\tilde{s}^{\prime}-\tilde{s}}}\log(c+\Delta\epsilon Y_{k})$
@@
    $Z_{u}:=\frac{\int_{\Sigma_{1}}\epsilon\exp H_{u}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)}{\int_{\Sigma_{1}}\exp H_{u}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)}.$
@@
    $\frac{\,\mathrm{d}}{\,\mathrm{d}u}Z_{u}=$ $\sum_{k\leq\Pi_{\tilde{s}}}(Y_{k}-X_{k})\bigg{(}\bigg{\langle}\frac{\epsilon}{c+\Delta\epsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg{\rangle}_{u}-\langle\epsilon\rangle_{u}\bigg{\langle}\frac{1}{c+\Delta\epsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg{\rangle}_{u}\bigg{)}$
    $\qquad+\sum_{x\in\Lambda((t^{\prime}-t)\tilde{\nu})}\big{(}\langle\epsilon\log(c+\Delta\epsilon x)\rangle_{u}-\langle\epsilon\rangle_{u}\langle\log(c+\Delta\epsilon x)\rangle_{u}\big{)}$
    $\qquad\qquad+\sum_{k\leq\Pi_{\tilde{s}^{\prime}-\tilde{s}}}\big{(}\langle\epsilon\log(c+\Delta\epsilon Y_{k})\rangle_{u}-\langle\epsilon\rangle_{u}\langle\log(c+\Delta\epsilon Y_{k})\rangle_{u}\big{)}.$
@@
    $\operatorname{\mathbb{E}}\lvert Z_{1}-Z_{0}\rvert$ $\leq C\bigg{(}\operatorname{\mathbb{E}}\sum_{k\leq\Pi_{\tilde{s}}}\lvert Y_{k}-X_{k}\rvert+\operatorname{\mathbb{E}}\lvert\Lambda((t^{\prime}-t)\tilde{\nu})\rvert+\operatorname{\mathbb{E}}\Pi_{\tilde{s}^{\prime}-\tilde{s}}\bigg{)}$
    $\leq C\big{(}\mu[-1,1]\operatorname{\mathbb{E}}\lvert Y_{1}-X_{1}\rvert+\lvert t^{\prime}-t\rvert+\lvert\mu^{\prime}[-1,1]-\mu[-1,1]\rvert\big{)}.$
@@
$W(\nu_{t,\mu},\nu_{t^{\prime},\mu^{\prime}})\leq C\big{(}\lvert t^{\prime}-t\rvert+\mu[-1,1]W\big{(}\overline{\mu},\overline{\mu}^{\prime}\big{)}+\lvert\mu[-1,1]-\mu^{\prime}[-1,1]\rvert\big{)}.$ (6.7)
@@
The Arzelà-Ascoli theorem gives a subsequence $(N_{k})_{k\geq 1}$ along which the free energy converges to some limit $f:\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}\to\mathbb{R}$. Arguing as in Proposition 4.3, find sequences $\smash{(N_{k}^{\pm},\lambda^{N_{k}^{\pm}})_{k\geq 1}}$ and measures $\nu^{\pm}\in\mathcal{M}_{p}$ such that the following holds. The sequences $\smash{(N_{k}^{\pm})_{k\geq 1}\subset(N_{k})_{k\geq 1}}$ increase to infinity, the multioverlap concentration (3.15) for the Hamiltonian (4.6) with perturbation parameters $\smash{(\lambda^{N_{k}^{\pm}})_{k\geq 1}}$ holds for all $n\geq 1$, the asymptotic spin array associated with the perturbation parameters $\smash{(\lambda^{N_{k}^{\pm}})_{k\geq 1}}$ is generated by $\nu^{\pm}$ according to (S1)-(S2), and
$\operatorname{\mathcal{P}}_{t,\mu}(\nu^{-})\leq f(t,\mu)\leq\operatorname{\mathcal{P}}_{t,\mu}(\nu^{+}).$ (6.8)
The idea is to now fix $C<+\infty$ as in the statement of Lemma 6.2, and show that for $t<C^{-1}$, the measure $\nu^{\pm}$ is equal to the unique fixed point $\nu_{t,\mu}\in\mathcal{M}_{p}$ of the operator $\Gamma_{t,\mu}$. This shows that the sequence of enriched free energies admits the unique subsequential limit $\operatorname{\mathcal{P}}_{t,\mu}(\nu_{t,\mu})$ for $t<C^{-1}$, and therefore converges to said limit in this small-time regime. It suffices to prove this at points $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathrm{Reg}(\mathcal{M}_{+})$ such that $f(t,\cdot)$ is Gateaux differentiable at $\mu$. Indeed, the equality $f(t,\mu)=\operatorname{\mathcal{P}}_{t,\mu}(\nu_{t,\mu})$ can then be extended to all measures $\mu\in\mathcal{M}_{+}$ using a density argument relying on Proposition 2.2 and the Lipschitz continuity results in Proposition 2.1 and Lemma 6.2. If $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathrm{Reg}(\mathcal{M}_{+})$ is such that $f(t,\cdot)$ is Gateaux differentiable at $\mu$ and $\nu\in\mathcal{M}_{p}$ is such that $G_{\nu}=D_{\mu}f(t,\mu,\cdot)$, then arguing as in the proof of Lemma 5.3 with $\nu^{*}$ replaced by $\nu^{\pm}$ and invoking Lemma 5.5, it is possible to show that
$G_{\nu^{\pm}}=G_{\nu}=D_{\mu}f(t,\mu,\cdot)=D_{\mu}\psi(\mu+t\nu,\cdot).$
By Lemma 5.2, the first of these equalities implies that $\nu^{\pm}=\nu$ which means that $\nu^{\pm}$ satisfies the fixed point equation
$G_{\nu^{\pm}}=D_{\mu}\psi(\mu+t\nu^{\pm},\cdot).$
Invoking Lemma 5.6 reveals that $\nu^{\pm}$ is a fixed point of the operator $\Gamma_{t,\mu}$, and is therefore equal to the unique fixed point $\nu_{t,\mu}\in\mathcal{M}_{p}$ whenever $t<C^{-1}$. Together with (6.8) this completes the proof. ∎

### v2

@@ 4 Cavity computation
Together with the previous display this gives the lower bound in (4.29). The upper bound in this inequality is proved in an identical manner by instead using that the $\limsup$ of a Cesàro sum is bounded from above by the $\limsup$ of its general term. This completes the proof. ∎
@@ 5 Critical point representation
To relate the Gateaux derivative of the limit free energy to that of the initial condition, it will be convenient to recall from [19, Remark 3.2] that for any $\mu\in\mathcal{M}_{+}$, the density of the Gateaux derivative $D_{\mu}\psi(\mu)$ is
$D_{\mu}\psi(\mu,x)=\operatorname{\mathbb{E}}\langle c+\Delta\varepsilon x\rangle_{*}\log\langle c+\Delta\varepsilon x\rangle_{*}-c-\Delta\overline{m}x,$ (5.1)
@@
$H^{\mu}(\varepsilon):=-\int_{-1}^{1}(c+\Delta\varepsilon x)\,\mathrm{d}\mu(x)+\sum_{x\in\Pi_{*}(\mu)}\log(c+\Delta\varepsilon x).$ (5.2)
Here $\Pi_{*}(\mu)$ denotes a Poisson point process on [1,1][-1,1] with mean measure $(c+\Delta\sigma^{*}x)\,\mathrm{d}\mu(x)$. This means that for any bounded and measurable function $f=f(\varepsilon^{1},\ldots,\varepsilon^{n})$ of finitely many replicas,
$\langle f(\varepsilon^{1},\ldots,\varepsilon^{n})\rangle_{*}:=\langle f\rangle_{*}:=\frac{\int_{\Sigma_{1}^{n}}f(\varepsilon^{1},\ldots,\varepsilon^{n})\prod_{\ell\leq n}\exp H^{\mu}(\varepsilon^{\ell})\,\mathrm{d}P^{*}(\varepsilon^{\ell})}{\big(\int_{\Sigma_{1}}\exp H^{\mu}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)\big)^{n}}.$ (5.3)
It will also be convenient to remember from the proof of [19, Lemma 2.4] that the Gateaux derivative of the enriched free energy (1.22) at $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ in the direction of a probability measure $\eta\in\Pr[-1,1]$ is given by
@@
where the bracket $\langle\cdot\rangle$ denotes the Gibbs average (2.14), the $(w_{i})_{i\geq 1}$ are i.i.d. random variables sampled from $\eta$, and $(\widetilde{G}_{i}^{w})_{i\leq N}$ are conditionally independent random variables taking values in $\{0,1\}$ with conditional distribution
@@
It follows that $\nu$ and $\nu^{\prime}$ have all the same moments of order $n\geq 2$, and are therefore equal by the Stone-Weierstrass theorem (see e.g. [20, Theorem A.10]). ∎
@@
$c-\frac{\Delta\operatorname{\mathbb{E}}w_{1}}{N+1}\sum_{i\leq N+1}\sigma_{i}+\operatorname{\operatorname{\mathcal{O}}}(N^{-1})=-c-\Delta\overline{m}\operatorname{\mathbb{E}}w_{1}+\operatorname{\operatorname{\mathcal{O}}}(N^{-1}).$
@@
$D_{\mu}\overline{F}_{N+1}(t,\mu,\lambda^{N};\eta)=(N+1)\operatorname{\mathbb{E}}\log\big\langle(c+\Delta\varepsilon w_{1})^{\widetilde{G}_{1}^{w}}\big\rangle_{N+1}-c-\Delta\overline{m}\operatorname{\mathbb{E}}w_{1}.$ (5.10)
@@
$D^{1}$ $:=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\big\langle(c+\Delta\varepsilon w_{1})^{\widetilde{G}_{1}^{w}}\exp z_{N}^{t,\mu}(\varepsilon,\sigma)\big\rangle^{\prime}-\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}^{t,\mu}(\varepsilon,\sigma)\rangle^{\prime}\,\mathrm{d}P^{*}(\varepsilon)$
@@
$D^{1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}e^{-c(t+s)-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}$ $(c+\Delta\varepsilon w_{1})^{{G^{\prime}}^{w}}\exp z_{N}^{\prime}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)$
$-\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}e^{-c(t+s)-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}\exp z_{N}^{\prime}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)+o_{N}(1),$
@@
$\mathbb{P}\big\{{G^{\prime}}^{w}=1\mid\varepsilon^{*},w_{1}\big\}:=\frac{c+\Delta\varepsilon^{*}w_{1}}{N+1}.$
@@
$D^{1}=\operatorname{\mathbb{E}}\log\big\langle(c+\Delta\varepsilon w_{1})^{{G^{\prime}}^{w}}\big\rangle_{*}+o_{N}(1)=\frac{1}{N+1}\operatorname{\mathbb{E}}\langle c+\Delta\varepsilon w_{1}\rangle_{*}\log\langle c+\Delta\varepsilon w_{1}\rangle_{*}+o_{N}(1).$
@@
$D_{\mu}f(t,\mu;\eta)=\operatorname{\mathbb{E}}\langle c+\Delta\varepsilon w_{1}\rangle_{*}\log\langle c+\Delta\varepsilon w_{1}\rangle_{*}-c-\Delta\overline{m}\operatorname{\mathbb{E}}w_{1}.$
@@
$D_{\mu}\psi(\mu+t\nu,x)=\operatorname{\mathbb{E}}g(\langle\varepsilon\rangle_{*}x)=(c+\Delta\operatorname{\mathbb{E}}\langle\varepsilon\rangle_{*}x)\log(c)+c\sum_{n\geq 2}\frac{(-\Delta/c)^{n}}{n(n-1)}\operatorname{\mathbb{E}}\langle\varepsilon\rangle_{*}^{n}x^{n}-c.$ (5.12)
The equation $G_{\nu}=D_{\mu}\psi(\mu+t\nu,\cdot)$ is therefore equivalent to the equality between all moments of $y_{1}$ and $\langle\varepsilon\rangle_{*}$ of order $2$ and higher. By the Stone-Weierstrass theorem, this is in turn equivalent to the equality between the laws of $y_{1}$ and $\langle\varepsilon\rangle_{*}$. These laws are equal if and only if the fixed point equation $\nu=\Gamma_{t,\mu}(\nu)$ is satisfied. This completes the proof. ∎
@@ 6 The disassortative and small-time settings
in such a way that $\varphi(1)=\overline{F}_{N}(t,\mu)$ and $\varphi(0)=\psi(\mu+t\nu)$. We denote by $x_{1}$ a random variable sampled from $\nu$. Recall from the derivative computations in [19, Corollaries 2.2 and 2.5] that
@@
[…]
@@
    Fix $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ as well as $\nu,\eta\in\mathcal{M}_{p}$. Conditionally on the ground truth signal $\varepsilon^{*}$ sampled from $P^{*}$, introduce the measures
    $\mathrm{d}\tilde{\nu}(x):=(c+\Delta\varepsilon^{*}x)\,\mathrm{d}\nu(x)\quad\text{and}\quad\,\mathrm{d}\tilde{\eta}(x):=(c+\Delta\varepsilon^{*}x)\,\mathrm{d}\eta(x).$ (6.5)
    We denote by $(X_{k})_{k\geq 1}$ and $(Y_{k})_{k\geq 1}$ i.i.d. samples from the probability measures $\nu^{\prime}$ and $\eta^{\prime}$ induced by $\tilde{\nu}$ and $\tilde{\eta}$, and let $\Pi_{t}$ be a Poisson random variable with mean $t(c+\Delta\varepsilon^{*}\overline{m})$. For each $u\in[0,1]$, we define the interpolating Hamiltonian $H_{u}$ on $\Sigma_{1}$ by
    $H_{u}(\varepsilon):=-\varepsilon\overline{m}+\sum_{x\in\Pi_{0,\mu}(0)}\log(c+\Delta\varepsilon x)+\sum_{k\leq\Pi_{t}}\log\big(c+\Delta\varepsilon(X_{k}+u(Y_{k}-X_{k}))\big)$
@@
    $Z_{u}:=\frac{\int_{\Sigma_{1}}\varepsilon\exp H_{u}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)}{\int_{\Sigma_{1}}\exp H_{u}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)}.$
@@
    $\frac{\,\mathrm{d}}{\,\mathrm{d}u}Z_{u}$ $=\sum_{k\leq\Pi_{t}}(Y_{k}-X_{k})\bigg(\bigg\langle\frac{\varepsilon}{c+\Delta\varepsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg\rangle_{u}-\langle\varepsilon\rangle_{u}\bigg\langle\frac{1}{c+\Delta\varepsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg\rangle_{u}\bigg).$
@@
    $F_{\nu^{\prime}}(x)=\frac{1}{c+\Delta\varepsilon^{*}\overline{m}}\bigg(cF_{\nu}(x)+\Delta\varepsilon^{*}\Big(xF_{\nu}(x)-\int_{-1}^{x}F_{\nu}(z)\,\mathrm{d}z\Big)\bigg).$
@@
    $H_{u}(\varepsilon):=-\varepsilon\overline{m}$ $+\sum_{x\in\Lambda(t\tilde{\nu})}\log(c+\Delta\varepsilon x)+u\sum_{x\in\Lambda((t^{\prime}-t)\tilde{\nu})}\log(c+\Delta\varepsilon x)$
    $+\sum_{k\leq\Pi_{\tilde{s}}}\log\big(c+\Delta\varepsilon(X_{k}+u(Y_{k}-X_{k}))\big)+u\sum_{k\leq\Pi_{\tilde{s}^{\prime}-\tilde{s}}}\log(c+\Delta\varepsilon Y_{k})$
@@
    $Z_{u}:=\frac{\int_{\Sigma_{1}}\varepsilon\exp H_{u}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)}{\int_{\Sigma_{1}}\exp H_{u}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)}.$
@@
    $\frac{\,\mathrm{d}}{\,\mathrm{d}u}Z_{u}=$ $\sum_{k\leq\Pi_{\tilde{s}}}(Y_{k}-X_{k})\bigg(\bigg\langle\frac{\varepsilon}{c+\Delta\varepsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg\rangle_{u}-\langle\varepsilon\rangle_{u}\bigg\langle\frac{1}{c+\Delta\varepsilon(X_{k}+u(Y_{k}-X_{k}))}\bigg\rangle_{u}\bigg)$
    $\qquad+\sum_{x\in\Lambda((t^{\prime}-t)\tilde{\nu})}\big(\langle\varepsilon\log(c+\Delta\varepsilon x)\rangle_{u}-\langle\varepsilon\rangle_{u}\langle\log(c+\Delta\varepsilon x)\rangle_{u}\big)$
    $\qquad\qquad+\sum_{k\leq\Pi_{\tilde{s}^{\prime}-\tilde{s}}}\big(\langle\varepsilon\log(c+\Delta\varepsilon Y_{k})\rangle_{u}-\langle\varepsilon\rangle_{u}\langle\log(c+\Delta\varepsilon Y_{k})\rangle_{u}\big).$
@@
[…]
@@
[…]
@@
The Arzelà-Ascoli theorem gives a subsequence $(N_{k})_{k\geq 1}$ along which the free energy converges to some limit $f:\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}\to\mathbb{R}$. With the notation $\nu_{t,\mu}$ as in Lemma 6.2, we first observe that if $f$ is Gateaux differentiable jointly in $(t,\mu)$ at some $(t,\mu)\in(0,C^{-1})\times\mathrm{Reg}(\mathcal{M}_{+})$, then one must have
$D_{\mu}f(t,\mu,\cdot)=G_{\nu_{t,\mu}}$ (6.8)
and
$\partial_{t}f(t,\mu)=\frac{1}{2}\int_{-1}^{1}G_{\nu_{t,\mu}}\,\,\mathrm{d}\nu_{t,\mu}.$ (6.9)
Indeed, the relation (6.8) follows from Lemmas 5.5, 5.6, and the uniqueness of the fixed point from Lemma 6.2, while (6.9) follows from the explicit representation of $\partial_{t}\overline{F}_{N}$ in (2.12), Lemmas 5.3, 5.5, and 5.6 which identify the limit law of the overlap array, and Proposition 2.3 which ensures the convergence of $\partial_{t}\overline{F}_{N_{k}}(t,\mu)$ to $\partial_{t}f(t,\mu)$.
Let us now fix $(t,\mu)\in(0,C^{-1})\times\mathcal{M}_{+}$, and for each $s\in\mathbb{R}$ sufficiently small, let $(t_{s},\mu_{s})\in\mathbb{R}_{>0}\times\mathrm{reg}(\mathcal{M}_{+})$ be a point of joint Gateaux differentiability of $f$ such that
$|t-t_{s}|+|F_{\mu}-F_{\mu_{s}}|_{L^{2}}\leq s^{2}.$
Such $(t_{s},\mu_{s})$ exists since the set of points of differentiability of $f$ is dense. By Proposition 2.1, we have that as $s>0$ tends to zero,
$s^{-1}(f(t+s,\mu)-f(t,\mu))$ $=s^{-1}(f(t_{s}+s,\mu_{s})-f(t_{s},\mu_{s}))+O(s)$
$\geq\frac{1}{2}\int_{-1}^{1}G_{\nu_{t_{s},\mu_{s}}}\,\,\mathrm{d}\nu_{t_{s},\mu_{s}}+O(s),$
where we used the convexity of $f$ in the $t$ variable from Proposition 2.9 and (6.9) to obtain the last line. The continuity of the mapping $(t,\mu)\mapsto\nu_{t,\mu}$ from Lemma 6.2 therefore implies that
$\liminf_{s\downarrow 0}s^{-1}(f(t+s,\mu)-f(t,\mu))\geq\frac{1}{2}\int_{-1}^{1}G_{\nu_{t,\mu}}\,\,\mathrm{d}\nu_{t,\mu}.$
In order to obtain the converse bound, we instead select $(t_{s}^{\prime},\mu_{s}^{\prime})\in\mathbb{R}_{>0}\times\mathrm{Reg}(\mathcal{M}_{+})$ such that $(t_{s}^{\prime}+s,\mu_{s}^{\prime})$ is a point of differentiability of $f$ and
$|t-t^{\prime}_{s}|+|F_{\mu}-F_{\mu^{\prime}_{s}}|_{L^{2}}\leq s^{2}.$
Using the Lipschitz continuity and the convexity of $f$ once more, we can then write
$s^{-1}(f(t+s,\mu)-f(t,\mu))$ $=s^{-1}(f(t^{\prime}_{s}+s,\mu^{\prime}_{s})-f(t^{\prime}_{s},\mu^{\prime}_{s}))+O(s)$
$\leq\frac{1}{2}\int_{-1}^{1}G_{\nu_{t_{s}^{\prime}+s,\mu^{\prime}_{s}}}\,\,\mathrm{d}\nu_{t_{s}^{\prime}+s,\mu^{\prime}_{s}}+O(s).$
Using again the continuity of the map $(t,\mu)\mapsto\nu_{t,\mu}$, we therefore conclude that
$\lim_{s\downarrow 0}s^{-1}(f(t+s,\mu)-f(t,\mu))=\frac{1}{2}\int_{-1}^{1}G_{\nu_{t,\mu}}\,\,\mathrm{d}\nu_{t,\mu}.$
The limit as $s$ tends to zero from the left can be handled similarly. We have thus shown that $f$ is differentiable in time everywhere on $(0,C^{-1})\times\mathcal{M}_{+}$, with derivative given by (6.9). Recalling also from [19, Lemma 3.1] that $f(0,\cdot)=\Psi$, we conclude that the function $f$ is determined uniquely as
$f(t,\mu)=\Psi(\mu)+\frac{1}{2}\int_{0}^{t}\int_{-1}^{1}G_{\nu_{s,\mu}}\,\,\mathrm{d}\nu_{s,\mu}\,\,\mathrm{d}s.$
In particular, the identity of $f$ does not depend on the subsequence, and we can thus conclude that $\overline{F}_{N}$ in fact converges along the full sequence. Proposition 1.3 then follows from Theorem 1.1 and the uniqueness of the fixed point on $(0,C^{-1})\times\mathcal{M}_{+}$. ∎